# NTSEBench: Cognitive Reasoning Benchmark for Vision Language Models

### v1

@@ Abstract
Cognitive textual and visual reasoning tasks, such as puzzles, series, and analogies, demand the ability to quickly reason, decipher, and evaluate patterns both textually and spatially. While LLMs and VLMs, through extensive training on large amounts of human-curated data, have attained a high level of pseudo-human intelligence in some common sense reasoning tasks, they still struggle with more complex reasoning tasks that require cognitive understanding. In this work, we introduce a new dataset, NTSEBench, designed to evaluate the cognitive multi-modal reasoning and problem-solving skills of large models. The dataset comprises 2,728 multiple-choice questions comprising of a total of 4,642 images across 26 categories sampled from the NTSE examination conducted nationwide in India, featuring both visual and textual general aptitude questions that do not rely on rote learning. We establish baselines on the dataset using state-of-the-art LLMs and VLMs. To facilitate a comparison between open source and propriety models, we propose four distinct modeling strategies to handle different modalities (text and images) in the dataset instances.
@@ 1 Introduction
Aptitude and reasoning tests have been essential for assessing intelligence and are considered strong indicators of problem-solving ability and abstract reasoning skills. Recent advancements in large language models (LLMs) have demonstrated their strong performance on IQ test questions, achieving high scores across five languages King (2023). These results suggest that LLMs are progressing towards some form of human intelligence at least in textual and language tasks.
@@
Figure 1: NTSEBench Examples: Shows three samples of textual, direction and spatial reasoning question from the proposed dataset. Solutions to these questions are not included here, but are provided in the dataset.
@@
The capabilities of LLMs models to perform on par with or exceed human baselines on various tasks—such as question answering, sentiment classification, text generation, visual question answering, coding challenges, data analytics, mathematical reasoning —continue to improve Srivastava et al. (2022); Bubeck et al. (2023). LLMs and VLMs have hence become the default benchmark for many zero or few shot text and vision-based tasks Brown et al. (2020); Wei et al. (2022). Training on huge amount of data from different domains has allowed LLMs to acquire human level performance in exams such as SAT, GRE, AP courses and on platforms leetcode Achiam et al. (2023).
@@
However, the results from King (2023) showed that LLMs performance is heavily skewed towards textual reasoning tasks like comprehension, textual analogies, opposites but do not perform good on other types of questions. These large models have achieved remarkable performance in many tests of human intelligence, they still fall short of human baselines in tasks requiring cognitive and logical thinking such as commonsense numerical & scientific reasoning, puzzles, and analogies Lu et al. (2023); Wang et al. (2023b). This challenge is evident not only in textual scenarios but also prominently in visual contexts. Most existing visual and multi-modal reasoning datasets in the literature are domain-specific, focusing on fields such as science, engineering, and medicine Yue et al. (2024); Zhang et al. (2024b); Sun et al. (2024) or they involve reasoning about real-world images. These datasets primarily assess tasks related to concrete scenarios or specific domains, which may not necessarily include spatial recognition, visual puzzle solving, abstract vision logic and pattern recognition. Visual puzzle tasks necessitate abstract and spatial reasoning abilities that are fundamentally different from those required for textual or language-based reasoning. However, none of the existing datasets explicitly focus on testing this aspect. In this study, we aim to address this research gap by introducing a novel benchmark dataset (NTSEBench) specifically crafted to evaluate the complex visual, textual, and multi-modal cognitive reasoning capabilities of large deep learning models such as LLMs and VLMs. Examples questions from the dataset are shown in Figure 1.
@@
NTSEBench is exclusively dedicated to establishing a benchmark for testing capabilities that do not rely on domain-specific knowledge or rote learning. Its primary contribution lies in evaluating the innate problem-solving skills inherent in human cognitive development and isolating where models lack by presenting well categorised data. NTSEBench comprises questions sourced from the nationwide talent search examination(NTSE) conducted in India by the directorate of education. It includes questions across various problem categories such as series, calendar & clock, direction sense, analogies, and more. These questions can be presented in text format, visual format, or both (multi-modal). We further evaluate the performance of recent LLMs and VLMs, including both Proprietary Achiam et al. (2023); Reid et al. (2024) and open source models Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024); Wang et al. (2023a); Bai et al. (2023). This work makes the following contributions:
@@
  • Set up bench mark baselines for multiple state of the art LLMs and VLMs, both open and proprietary source models.
  • Evaluate performance with distinct modelling strategies to handle multi-modal input handling for reasoning questions.
2 NTSEBench
The National Talent Search Examination (NTSE), administered by the National Council of Educational Research and Training (NCERT) in India since 1963, is a nationwide exam for 10th-grade students. The exam consists of two sections designed to assess a wide range of analytical skills: the Mental Ability Test (MAT) and the Scholastic Aptitude Test (SAT). The MAT section evaluates students’ general aptitude, critical thinking, logical and spatial reasoning, and analytical problem-solving skills (for both textual and visual problems). In contrast, the SAT assesses their domain-specific knowledge in science and mathematics. All questions in the NTSE are multiple-choice (MCQ) with one correct option out of four. A set of questions may have a common direction, which could be common instructions/information about the problems. Questions and options can be text/image or a combination of both , i.e., multi modal. Students are given 2 hours to solve 100 questions for each section. We hence aim to create a dataset focused on cognitive reasoning abilities or MAT-type questions and developed a benchmark dataset called NTSEBench to evaluate LLMs and VLMs.
@@ 2.1 Dataset Sources
To create the dataset, we primarily utilized previous years’ NTSE papers publicly released along with their solutions provided by Resonance  [^1] Paper Links, a coaching institute in India that offers paid lessons to students to help them improve their NTSE scores. Additionally, we used NTSE preparation materials, such as a reference book titled A Modern Approach to Verbal and Non-Verbal Reasoning  [^2] A Modern Approach to Verbal and Non-Verbal Reasoning which includes additional logical reasoning problems. We also incorporated content from the another book titled Study Guide for NTSE  [^3] Study Guide for NTSE to construct our dataset.
2.2 Extraction Pipeline
As outlined in the dataset sources, we used previous years’ NTSE papers to create the dataset. We create a pipeline and extract questions from these sources with human intervention to monitor mistakes.
The data extraction pipeline involves first processing the PDF through MathPix OCR [^4] MathPix OCR to generate a Word file, which was then manually corrected for any errors. Next, we used the docxlatex 555docxlatex library to convert all equations into LaTeX expressions. Finally, we leveraged the PyMuPDF  [^6] PyMuPDF library to extract all text and images, extracting (1) Textual data i.e. Direction(Extra guidance on the context of the question), the Question itself,the correct Answer option and the Solution alongside (2) Vision Based data i.e. relevant images which we segregate into Direction images, Problem images, Option images, and Solution images.
A total of 2,728 MCQ (multiple-choice questions) comprising of a total of 4,642 images across 26 categories questions are extracted and NTSEBench is created along with the necessary metadata.
2.3 Dataset Details
Problem Categories: NTSEBench encompasses several broad problem categories, each designed to test a distinct set of skills. Questions from these categories frequently appear in NTSE exams year after year. The description of various categories is as in Appendix Table 5.
Problem Sub categories The above categories are further subdivided based on the different kind of reasoning required. Table 1 lists each sub-category for Text Only questions and Text+Vision questions, along with the respective count for each category. As in Table 1 most categories are represented well in the dataset with mean number of questions being 104 across categories.
@@ 2.3 Dataset Details
Table 1: Problem categories and number of samples
Modality Variations Since NTSEBench has multi-modal questions, options and solutions, we have results in eight different combination of modality type that can occur for question-options-solution triplet. Table 2 shows the count of each triplet option.
Question Options Solutions # Samples
1199
381
70
18
330
126
403
201
Table 2: Modality Variations Question Count: Tick(✓) mark indicates whether question, option or solution contains image.
3 Benchmarking using LLMs and VLMs
@@
3.1 Modelling Strategies
To accommodate diverse modalities for question, option, and solution triplets, a single model supporting all combinations can be created by converting all images to text description via caption generation Li et al. (2022, 2023b) or convert all text to images and rely on OCR capabilities of the modelFujitake (2024); Zhao et al. (2023); Shi et al. (2023) to under and answer the question correctly. Both of these modelling accuracy depends on other sub-tasks, i.e., caption generation and OCR.
When questions and options incorporate images, we move into the domain of utilizing multiple-image inputs for model reasoning. This functionality is supported through APIs in models such as GPT-4o and GeminiReid et al. (2024)(proprietary). Due to these sub-task and input dependencies, in order to assess the reasoning capabilities of the underlying model and achieve a fair and comprehensive comparison of baseline models, we propose using four distinct modeling strategies.
3.1.1 Standard QA Model for Text-only questions
@@
For instances where question type($J$) for questions($Q$), options($O$) and solutions($S$) is text($T$), we use a standard text-based QA model like GPT3.5-Turbo or Llama3-70bAI@Meta (2024) or Mixtral8x7bJiang et al. (2024).
3.1.2 Image-Only Model for all Questions
With increasing OCR capabilities of several models Shi et al. (2023); Fujitake (2024); Zhao et al. (2023) for both document and hand writing recognition, especially for printed and scanned text.
We propose a modelling approach where all Questions and Options are presented to the model as a single image. The single image is created by combining all necessary textual and visual content as it appears in the examination paper, i.e., a snapshot of the question and options be it textual or visual question.
The model is presented with a single image with a system prompt directing the model to generate the result in a specified format. The advantage of this modelling approach is that it can be applied for all modality types of $D_{i}$.
3.1.3 Interleaved model for Text with Multiple Images
@@
In this approach, we integrate text with multiple images to create an interwoven context. This method involves placing related textual and visual elements in close proximity, enhancing the model’s ability to draw connections between them.
3.1.4 Standard VQA model with Single Stitched Image and Annotated Text
@@
Open-source models do not support the integration of text and images within a single prompt. To enable a fair comparison, we propose an alternative modeling strategy where all question and option images are stitched into a single image, labeled as Figure 1, Figure 2, etc., and accompanied by a textual prompt describing different parts of the image. The prompt refers to the images in a structured manner, guiding the model’s attention to relevant visual details. This composite image is then used to test the model’s performance, assessing its ability to understand and respond to questions based on the integrated visual and textual information.
@@
Example inputs for each of the above modelling strategy proposed are shown in Figure 2.
3.2 Prompting Strategies
We mainly employed two main modelling stragegies for setting up all the baselines on the proposed dataset, (A) Zero Shot and (B) Few shot, both using Chain of Thought Wei et al. (2022) prompting to get better results.
Zero Shot COT: Model is given prompt containing question and corresponding options and asked to choose the right option. The model is also asked to provide an reasoning behind the choice made by the model. Few Shot COT: In few shot chain of thought (COT), few samples($N$) with question, options and solution triplets($D_{i}$) are added to the prompts before asking the question under test. $N$, the number of exemplars are chosen keep in mind the token length supported by the model.
3.3 Implementation Details
@@
We evaluate NTSEBench using multiple open and closed source LLMs Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023) and VLMs Reid et al. (2024); Bai et al. (2023); Dong et al. (2024); Wang et al. (2023a) . We used low temperature setting to promote reproducibility.
LLMs: GPT-3.5-Turbo, Llama3-70bAI@Meta (2024), Mixtral8x7bJiang et al. (2024) using the Standard QA strategy with both zero shot COT and few shot COT.
Open-Source VLMs [^7] Evaluated on A6000 machine with half precision used for internLM: QWEN-VL-chat-7b Bai et al. (2023), CogVLM-2-Llama3-chat-19B Wang et al. (2023a), internlm-xcomposer2-vl-7b Dong et al. (2024) using the Standard VQA and Image-Only strategies with zero shot COT.  [^8] Refer to technical discussion section for details on why few shot COT is challenging with open source models.
Proprietary VLMs: Gemini-1.5-Pro Reid et al. (2024), GPT-4o have been evaluated with Interleaving, Standard VQA and Image-Only approaches with both zero-shot COT and few-shot prompting COT. We provide the prompts and the hyper-parameters in the appendix A.2 and . We have also evaluated the cheaper and faster version of Gemini, namely Gemini-1.5-Flash, which has shown comparable performance to Gemini-1.5-Pro.
Answer Extraction and Evaluation.
The answer extraction module is built using a rule-based approach to identify the correct option from the generated response, handling cases where the requested format is not followed Gupta et al. (2023). We used percentage accuracy as our metric for evaluation across all the models. If the model chooses the right option in the MCQ question label is set of True, else label is False. There is no partial marking for partially correct answers.
@@ 4 Results and Analysis
  • How well do large deep learning models perform on advanced textual, visual, and multi-modal reasoning questions? How challenging is cognitive reasoning for the current state-of-the-art models?
  • Are proprietary models superior to open-source models, and by what margin? Are there specific categories where proprietary models significantly outperform open-source models?
  • Do different modeling strategies affect a model’s accuracy? Does OCR impact the reasoning accuracy of models?
4.1 Results
Results for text only questions using Standard QA and Image Only(implicit OCR) modelling strategy are shown in Table 3. For Standard QA strategy we report results for zero shot setting in all eight models listed in the section above. For few shot setting, three models are excluded(CogVLM-2, Qwen-VL-Chat, InternLM-XComposer2) from the results. For Image Only modelling results are reported for five models which suport vision or multi-modal inputs for zero shot results. For some categories the solution contains Columns in the table refers to different types of questions such as SER (Series), ALP (Alphabet Test), ODO (Odd one out), ANA (Analogy), COD (Coding-Decoding), NUM (Number and Ranking), BLR (Blood Relation), MTO (Mathematical Operations), PUZ (Puzzle Test), SYL (Syllogisms), STC (Statement & Conclusions) and DAT (Data Sufficiency).
@@ 4.1 Results
Table 3: Text Only Results: Zero-shot and Few-shot performance of different models across various Text-Only categories. We report result using two different modelling strategies Image Only and Standard QA. italics font for propriety models, i.e., money or API access is required to run these models. The # is due to the category’s solution contains thus restricting few shot on text-only models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
Results for multi-modality question are reported in Table 4. Results are reported using three different modelling strategies listed in the section above namely Interleaved, Image Only and Standard VQA. Column in the table refer to different categories of visual text reasoning questions in NTSEBench such as DIR (Direction Sense), VEN (Venn Diagrams), TIM (Time and Clock), MIS (Missing Character), NVS (Non-Verbal Series), NVO (Non-Verbal odd one out), NVA (Non-Verbal Analogy), INC (Incomplete Figure), MIR (Mirror, Water and Images), CUB (Cube and Dice), PAP (Paper Folding & Cutting), EMB (Embedded Figure), FIG (Figure Partition), DOT (Dot Problem).
Proprietary models outperform Open Source models. From results in table 3 and 4, we can clearly observe that Proprietary models such as Gemini Pro 1.5 and GPT-4o , outperform other open source models in nearly every category question category. Proprietary models are generally twice as accurate as open-source models on NTSEBench questions, including both text-based and multi-modal questions, across all the different modeling strategies proposed in the section above. Results from Gemini Pro 1.5 outperform GPT-4o for the majority of modeling strategies on both text-based and multi-modal questions. Proprietary models undergo regular updates with refined weights tailored to specific tasks, as discussed in the literature Reid et al. (2024); Achiam et al. (2023). Consequently, we posit that proprietary models fine-tuned on domain-specific reasoning datasets Lu et al. (2023); He et al. (2024) have acquired advanced reasoning capabilities that contribute to their enhanced accuracy.
LLAMA-3-70B is the best-performing open-source model for text-only questions, while Qwen-VL-Chat prevails as the top open-source VLM model for multi-modal questions.
Modelling Strategy is important. For text-only questions, the Standard QA strategy clearly outperforms Image Only modeling. Introducing OCR on top of reasoning tends to confuse models or exacerbate the difficulty of the task. This effect is particularly noticeable with smaller open-source models, which struggle to accurately extract characters and integrate them into context. However, proprietary models like GPT-4o and Gemini-Pro still achieve superior results using Image Only compared to open-source models employing Standard QA or text-only processing alone.
Few-shot results present a mixed picture: while models like Mixtral and GPT-4o show improved performance with added exemplars, others experience a significant decline. We observe from tables 3 and 4 that proprietary models generally exhibit a smaller decline in performance compared to open-source models in such scenarios.
Interleaving text and images performs better than Standard VQA and Image Only strategy for most categories. This highlights the importance of presenting text and images separately and in a more fine-grained manner, providing appropriate context, and integrating the image with text as distinct entities. This also highlights that decomposing the problems help. Our results in Table 4 show that such an approach can significantly enhance and improve the performance of Vision-and-Language Models (VLMs). GeminiPro Reid et al. (2024) achieves the highest accuracy, highlighting the advanced capabilities of Gemini models in reasoning about different parts of images.
From table 4, we can clearly see that few-shot prompting consistently performs poorly for all VLM models compared to zero-shot prompting. This suggests that replicating or inferring underlying reasoning from multi-modal data is notably challenging.
Multi-modal reasoning is significantly harder. Comparing the best and worst performing models in Tables 3 and 4, it’s evident that multi-modal reasoning is considerably more challenging than textual reasoning for current state-of-the-art VLMs. The highest accuracy on multi-modal questions is below 35%, whereas for textual reasoning, the best-performing model achieves an accuracy of over 62%. This underscores the significant gap in reasoning capabilities between state-of-the-art VLMs and their LLM counterparts, highlighting the necessity for improved architectures and datasets tailored specifically for VLM models.
@@
Zero Shot CogVLM-2 18.75 18.02 25.49 14.96 18.95 20 8 12.77 7.61 19.1 16.67 12.5 12.68 4.35 14.98
Qwen-VL-Chat 21.05 26.13 27.45 22.22 26.32 21.43 17 21.28 19.57 25.84 25 18.75 18.31 17.39 21.98
InternLM-XComposer2 20.83 20.72 15.69 17.32 15.79 11.43 10 14.89 8.7 19.1 10.42 11.46 22.54 8.7 14.82

### v2

@@ Abstract
Cognitive textual and visual reasoning tasks, including puzzles, series, and analogies, demand the ability to quickly reason, decipher, and evaluate patterns both textually and spatially. Due to extensive training on vast amounts of human-curated data, LLMs and VLMs excel in common-sense reasoning tasks, however still struggle with more complex reasoning that demands deeper cognitive understanding. We introduce NTSEBench, a new dataset designed to evaluate cognitive multi-modal reasoning and problem-solving skills of large models. The dataset contains 2,728 multiple-choice questions, accompanied by a total of 4,642 images, categorized into 26 different types. These questions are drawn from the nationwide NTSE examination in India and feature a mix of visual and textual general aptitude challenges, designed to assess intelligence and critical thinking skills beyond mere rote learning. We establish baselines on the dataset using state-of-the-art LLMs and VLMs. To facilitate a comparison between open source and propriety models, we propose four distinct modeling strategies to handle different modalities—text and images—in the dataset instances.
@@ 1 Introduction
Aptitude and reasoning tests have been essential for assessing intelligence and are considered strong indicators of problem-solving ability and abstract reasoning skills Stern (1914). Recent advancements in large language models (LLMs) have demonstrated their strong performance on IQ test questions, achieving high scores across many languages King (2023). These results indicate that LLMs are advancing toward pseudo human-like intelligence, particularly in text and language tasks.
@@
Figure 1: NTSEBench Examples: Shows three samples of textual, direction, and spatial reasoning questions from the proposed dataset. Solutions to these questions are not included here but are provided in the dataset.
@@
The capabilities of LLM models rivals humans on various tasks—question answering(QA), sentiment classification, text generation, visual QA, coding challenges, and mathematical reasoning, to name a few Srivastava et al. (2022); Bubeck et al. (2023). LLMs and Vision Language Models (VLMs) also know as Multimodal Large Language Models(MLLM) have hence become the default benchmark for many zero or few-shot text and vision-based tasks Brown et al. (2020); Wei et al. (2022). Training on vast datasets from diverse domains has enabled LLMs to achieve human-level performance in SAT, GRE, and AP exams, and on platforms like LeetCode Achiam et al. (2023); Touvron et al. (2023); Dubey et al. (2024).
@@
The results from King (2023) indicate that LLMs excel in tasks involving textual reasoning, such as comprehension, analogies, and identifying opposites, but struggle with other types of questions. While Multi-modal Large Language models have demonstrated remarkable performance in many tests of human intelligence, yet they still fall short of human baselines in tasks that require critical and logical thinking, such as commonsense-, numerical- and scientific-reasoning, puzzles, and analogies Most existing visual and multi-modal reasoning datasets are domain-specific, concentrating on fields like science, engineering, and medicine Yue et al. (2024); Zhang et al. (2024b); Sun et al. (2024). These datasets primarily focus on tasks related to concrete scenarios or specific domains, often requiring domain knowledge and rote learning for high performance. However, they do not adequately assess intelligence as a function of cognitive/critical reasoning skills such as spatial recognition, visual puzzle solving, abstract reasoning, or pattern recognition. In this study, we aim to address this research gap by introducing a novel benchmark dataset, NTSEBench, specifically created to evaluate the complex visual, textual, and multi-modal cognitive reasoning capabilities of large deep learning models such as LLMs and VLMs. Given the potential of LLM/VLMs to replace human jobs in the future, our dataset aims to answer the following question: Do LLM/VLMs possess some form of general human intelligence?. Examples of questions from the proposed dataset are shown in Figure 1.
@@
NTSEBench is dedicated to establishing a benchmark for testing capabilities that do not rely on domain-specific knowledge or rote learning. Its primary contribution lies in evaluating the innate problem-solving skills inherent in human cognitive development and isolating where models lack by presenting well-categorized data. It comprises of questions sourced from the nationwide talent search examination (NTSE) conducted in India. These questions can be presented in text format, visual format, or both (multi-modal). We evaluate the performance of recent LLMs and VLMs, including both Proprietary Achiam et al. (2023); Reid et al. (2024) and open source models Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024a); Wang et al. (2023a); Bai et al. (2023); Li et al. (2024); Lu et al. (2024) on our dataset and perform in-depth error analysis of LLM responses to pinpoint areas of weakness and evaluate their overall robustness. Our work makes the following contributions:
@@
  • Establish baselines using a diverse range of state-of-the-art LLMs and VLMs on the proposed dataset, incorporating both open-source and proprietary models.
  • Assess performance using various modeling strategies to effectively handle multimodal (multi image inputs as well) inputs for reasoning-based questions.
2 NTSEBench Benchmark
NTSE Exam. The National Talent Search Examination (NTSE), administered by the National Council of Educational Research and Training (NCERT) in India since 1963, is a nationwide exam for secondary-grade students. The exam consists of two sections designed to assess a wide range of analytical skills: the Mental Ability Test (MAT) and the Scholastic Aptitude Test (SAT). The MAT section evaluates students’ general aptitude, critical thinking, logical and spatial reasoning, and analytical problem-solving skills (for both textual and visual problems). In contrast, the SAT assesses their domain-specific knowledge in science and mathematics. All questions in the NTSE are multiple-choice (MCQ) with one correct option. Questions and options can be text/image or a combination of both, i.e., multi-modal. We aim to create a dataset focused on cognitive reasoning abilities (MAT-type questions).
Cognitive Reasoning. Cognitive understanding in the context of NTSEBench refers to the ability to process information, recognize patterns, draw inferences, and solve problems using critical, logical and analytical reasoning. This aligns with fundamental concepts in cognitive science, such as problem-solving, pattern recognition, and inferential reasoning Wang and Chiew (2010). It encompasses advanced reasoning skills typically found in a small subset of the population, generally individuals with very high IQs. To emphasize this distinction, we use the term cognitive reasoning for our dataset, differentiating it from common sense reasoning tasks Sakaguchi et al. (2021); Talmor et al. (2019). NTSEBench assesses these reasoning abilities through diverse question categories, each targeting a different cognitive dimension:
  • Pattern Recognition: Categories like Series (Numerical, Alphabetical, Alphanumeric), Missing Character, Non-Verbal Series, and Dot Problem test the ability to identify and extend patterns, which is crucial for understanding sequences and predicting outcomes.
  • Logical Deduction: Blood Relation, Syllogisms, Statement & Conclusions, and Data Sufficiency categories focus on making inferences and drawing conclusions based on the given information, reflecting the core of logical reasoning.
  • Spatial Reasoning: Direction Sense, Cube and Dice, Paper Folding & Cutting, and Embedded Figure assess the ability to visualize and manipulate objects in space, which is essential for understanding spatial relationships.
  • Relational Reasoning: Analogy and Non-Verbal Analogy categories evaluate the understanding of relationships between items and the ability to transfer this understanding to new contexts, a key component of relational reasoning.
  • Quantitative Analysis: Number and Ranking, Mathematical Operations, Time and Clock, and Figure Partition test numerical problem-solving skills and the ability to manage quantitative data effectively.
  • Classification and Categorization: Classification/Odd One Out, Non-Verbal Classification/Odd One Out categories measure the ability to group items based on shared attributes and identify outliers, highlighting skills in distinguishing unique characteristics and grouping.
  • Contextual Interpretation: Looking specific details, instructions, or constraints that are critical to understanding and solving the problem.
  • Verbal Reasoning: Understanding semantic relationships, word meanings, and analogies.
Unlike other benchmarks that focus on specific academic domains or individual cognitive dimensions as stated above, NTSEBench emphasizes a wide range of cognitive skills, offering a more comprehensive assessment of reasoning abilities. We have expanded on our proposed dataset’s global relevance in Appendix  A.1.
@@ 2.1 Dataset Sources
We created the dataset using past NTSE papers and solutions from Resonance. Additionally, we used NTSE preparation materials, such as a reference book titled A Modern Approach to Verbal and Non-Verbal Reasoning which includes additional logical reasoning problems. We also incorporated content from another book titled Study Guide for NTSE to construct our dataset. The question extraction process is detailed in Appendix A.2.1. The full dataset will be released upon acceptance, with example questions in Figure 1 and Appendix B.
Problem Categories: NTSEBench encompasses several problem categories, each designed to test a distinct set of skills. Questions from these categories frequently appear in NTSE exams year after year. A detailed description of each category is done in Appendix Table 4 and examples shown in Appendix section. More dataset related information is present in Appendix section A.2.
@@
Table 1: NTSEBench categories count: Problem categories with different input modality types and number of samples for each.
3 Models: LLMs and VLMs
@@
Modelling Strategies.
Evaluating the reasoning abilities of large language models (LLMs) with text-based questions is straightforward. For vision-language models (VLMs), reasoning with vision-text questions is generally straightforward. Some API access model models, like GPT-4o and Gemini, support multi-image inputs, while many others do not (open-source models like LLaVA-OneVision and Ovis are emerging with this capability). To address these task-specific and input-related dependencies, we propose four strategies to fairly evaluate the reasoning abilities of both open-source and API-based models.
@@
  • Standard QA. For instances where question type($J$) for questions($Q$), options($O$) and solutions($S$) is text($T$), we use a standard text-based QA model like GPT3.5-Turbo or Llama3-70b AI@Meta (2024) or Mixtral8x7b Jiang et al. (2024a).
  • Image-Only. We propose a modeling approach where questions and all the options are presented to the model as a single image. This image consolidates all relevant textual and visual content exactly as it appears in the examination paper, effectively capturing the entire question, including both textual and visuals. This strategy utilizes the OCR capabilities of VLM models to interpret and analyze the content, enabling them to process both text and visual elements within the same input Shi et al. (2023); Fujitake (2024); Zhao et al. (2023). The key advantage of this approach is its applicability across all modality types.
@@
  • Interleaved model. In this approach, we integrate text with multiple images to create an interwoven context. This method involves placing related textual and visual elements in proximity, enhancing the model’s ability to draw connections between them.
@@
  • Standard VQA. Open-source models typically lack the capability to integrate text and images within a single prompt. To enable fair comparisons, we propose an alternative modeling strategy where the question and option images are combined into a single composite image, labeled as Figure 1, Figure 2, etc. This composite image is accompanied by a structured textual prompt that describes different parts of the image, directing the model’s attention to relevant visual details. The composite image and prompt are then used to evaluate the model’s performance, testing its ability to interpret and respond to questions based on the integrated visual and textual information.
@@
Example inputs for each of the above modelling strategy proposed are shown in Appendix Figure 3.
Prompting Strategies. We mainly employed two main prompting strategies for setting up all the baselines on the proposed dataset: (A) Zero Shot COT: The model is presented with a prompt that includes a question and a set of answer options. It is tasked with selecting the correct option and providing an explanation that justifies its choice. (B) Few Shot COT: In few-shot chain-of-thought (COT) prompting, a set of $N$ exemplar triplets—each containing a question, options, and a solution ($D_{i}$)—is included in the prompt before presenting the test question. The number of exemplars $N$ is selected based on the token limit supported by the model.
@@
Implementation Details. We evaluate NTSEBench using multiple open and closed source LLMs Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023) and VLMs Reid et al. (2024); Bai et al. (2023); Dong et al. (2024); Wang et al. (2023a). We used low temperature setting to promote reproducibility. Details on models and hyperparameter can be found in Appendix  A.5.
In the few-shot settings, the questions were sorted by solution length, then an annotator picked the three examples with the most comprehensive responses.
Answer Extraction and Evaluation. The answer extraction module uses a rule-based approach to identify the correct option from the response Gupta et al. (2023). Responses where the answers can’t be extracted by module are evaluated by humans. We used percentage accuracy as our metric for evaluation across all the models. If the model chooses the right option in the MCQ question, label is set of True, else label is False. Option Bias Ablation. We conducted option shuffling experiments to assess whether model performance is influenced by the position of the correct option and to detect potential bias. Four dataset variations were created, placing the correct answer in positions 1, 2, 3, and 4, respectively. Each experiment was run in three rounds, and results were averaged to ensure robustness.
@@ 4 Results and Analysis
1. How well do LLMs and VLMs perform on advanced textual, visual, and multi-modal reasoning questions? How challenging is cognitive reasoning for the current state-of-the-art models?
2. Are proprietary models superior to open-source models, and by what margin? Are there specific categories where proprietary models significantly outperform open-source models?
3. Do different modeling strategies affect a model’s accuracy? Does OCR impact the reasoning accuracy of models?
4. Does shuffling the options in questions impact model performance, indicating a potential option bias in large models?
Results for text only questions using Standard QA and Image Only (implicit OCR) modelling strategy are shown in Table 2. For Standard QA strategy, we report results for zero shot setting in all 10 models listed in the appendix  A.5. For Image Only modelling results are reported for five models which support vision or multi-modal inputs for zero shot results. Columns in the table refers to different types of questions in NTSEBench such as SER (series), ALP (alphabet Test), ODO (odd one out), ANA (analogy), COD (coding-decoding), NUM (number and ranking), BLR (blood relation), MTO (mathematical operations), PUZ (puzzle test), SYL (syllogisms), STC (statement & conclusions) and DAT (data sufficiency).
@@
Advanced Reasoning Models
OpenAI o1-preview 80.62 90.22 84.05 73.13 85 85.83 83.61 83.7 84.81 81.08 72.28 83.33 81.88
Table 2: Text Only Question Results. Zero-shot and Few-shot performance of different models across various text-only categories. We report results using two different modelling strategies Image Only and Standard QA. italics font for propriety models, i.e., money or API access is required to run these models. The # is due to the category’s solution contains images thus restricting few shot on text-only models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
@@
Zero Shot CogVLM-2 18.75 18.02 25.49 14.96 18.95 20 8.00 12.77 7.61 19.10 16.67 12.50 12.68 4.35 14.98
Qwen-VL-Chat 21.05 26.13 27.45 22.22 26.32 21.43 17.00 21.28 19.57 25.84 25 18.75 18.31 17.39 21.98
InternLM-XComposer2 20.83 20.72 15.69 17.32 15.79 11.43 10.00 14.89 8.70 19.10 10.42 11.46 22.54 8.70 14.82